# Strong coupling non-Markovian quantum thermodynamics using exact dynamics of a central spin surrounded by a spin bath and the role of a canonical Hamiltonian

### v1

@@ III.1 The canonical Hamiltonian
$\Im(\cdot)$ denotes the imaginary part of ()($\cdot$)( ⋅ ). Further, $D(0)=\frac{\omega_{0}}{2}$.
@@
In Fig. 1, we plot the factor $D(t)$ with time $t$. As time progresses, the factor $D(t)$ shows an oscillatory behavior with time around the line of $\frac{\omega_{0}}{2}$; however, asymmetric. The minimal dissipator can be found using Eq. (II).
@@
IV Quantum thermodynamical properties of the central spin model
Using the exact dynamics of the central spin system discussed above, we calculate the entropy production, heat, and work. Further, we calculate the maximum amount of work extractable from the system; that is, its ergotropy.
IV.1 Entropy production, heat and work
Considering the global system-bath dynamics for arbitrary initial states of the system and the bath [Eq. (12)], a formulation for the entropy production was developed in [86, 28]. Here, the dynamical evolution of the system is obtained by tracing out the bath, which is the origin of the irreversibility. The entropy production accounts for the contribution of two processes, one where we discard any information stored locally in the state of the bath, and the other is the non-local information shared between the system and the bath. The entropy production is given by
Further, when we consider the thermal state to be the initial state of the bath, which is the case considered here [Eq. (14)], the entropy production boils down to
$\Sigma=\Delta S_{S}+\beta Q_{B},$ (28)
is the total change in energy of the environment during the unitary evolution of the total system, where $\rho_{B}(t)={\rm Tr}_{S}[\rho_{SB}(t)]$. Equation (28) with $\Sigma\geq 0$ is the standard form of the second law of thermodynamics.
Furthermore, the change in the internal energy of the central spin system is given by
$\Delta U_{S}=W-Q_{B},$ (31)
$z(t)$ $=2\left[\alpha(t)\rho_{00}(0)+\eta(t)\rho_{11}(0)\right]-1,$
$x(t)$ $=2\Re\left\{\delta(t)\rho_{01}(0)\right\},$
$y(t)$ $=-2\Im\left\{\delta(t)\rho_{01}(0)\right\}.$ (33)
$\Delta U_{S}=\frac{\omega_{0}}{2}\left[z(t)-z(0)\right].$ (34)
$\rho_{B}(t)={\rm Tr}_{S}\left\{e^{-iHt}\rho_{S}(0)\otimes\rho_{B}(0)e^{iHt}\right\}.$ (35)
[…]
$W$ $={\rm Tr}\left[V\left\{\rho_{SB}(0)-\rho_{SB}(t)\right\}\right].$ (39)
This form is particularly useful as it brings out how, in a general scenario of strong system-bath coupling and under constant Hamiltonians, the work done by the bath can be explicitly computed, saving the need to calculate the change in the internal energies of the system and the bath. It is also worth mentioning that this amount of work must be supplied to turn off the interaction between the system and the bath. Interestingly, from Fig. 2(c), we observe that the amount of work required to turn off the interaction is zero in the case of weak coupling ($\epsilon=0.01$) and increases with an increase in the coupling strength. This observation is consistent with the literature [65].
The variations of the above-mentioned thermodynamic quantities with time are plotted in Fig. 2. We observe that the change in the internal energy of the system $\Delta U_{S}$ is positive and undergoes revivals in time, as depicted by the oscillations. For higher values of system-bath interaction parameter $\epsilon$, we observe well-defined wavepacket-like structures. Further, at a very low value of $\epsilon$, the system’s internal energy change is approximately zero. The above observation matches the variations in the change in the energy of the bath $Q_{B}$ and the work done by the bath $W$. The values of the $Q_{B}$ are zero for weak coupling ($\epsilon=0.01$), which is consistent with the GKLS condition, and are otherwise negative, denoting that the energy of the bath decreases during the dynamics, and the positive values of $W$ denote that the bath works on the system. This observation is consistent with the first law of thermodynamics; that is, the energy of the bath decreases while it works on the system. In the process, the internal energy of the system increases.
IV.2 Ergotropy
Here, $H_{S}$ is the system Hamiltonian, and $\rho_{S}^{p}$ is a final state that evolved from a state potent of work production $\rho_{S}(t)$, that is, the passive state. After an external time-dependent cyclic potential has done the maximum work, the state has turned into a passive state, which would be, in general, different from the Gibbs state attained for macroscopic systems. Once the passive state is reached, no more work can be extracted from it through any cyclic process. The spectral decomposition of the passive state and the Hamiltonian of the system is given by
$\rho_{S}^{p}=\sum_{j}r_{j}\ket{\epsilon_{j}}\bra{\epsilon_{j}},$
$H_{S}=\sum_{k}\epsilon_{k}\ket{\epsilon_{k}}\bra{\epsilon_{k}},$ (41)
V Role of the canonical Hamiltonian
In the preceding section, we discussed the quantum thermodynamics of the central spin system using the system’s bare Hamiltonian $H_{S}$ and the reduced state of the system at any time $t$. However, the dynamics of the reduced state of the system is obtained from a quantum master equation having a canonical Hamiltonian $H^{\rm can}_{S}$, as discussed in Sec. III.1. Therefore, it becomes natural to discuss the role of the canonical Hamiltonian in calculating the thermodynamical observables [68, 69]. To this end, we calculate the change in energy of the system by replacing the system’s Hamiltonian with the canonical Hamiltonian, that is,
@@ V Role of the canonical Hamiltonian
One may argue that the canonical Hamiltonian contains the effect of the bath even when the bath is partially traced and that it should be used to calculate expectation values of the thermodynamical quantities. By this argument, we observe that the system’s Hamiltonian becomes time-dependent even if we take a time-independent initial Hamiltonian of the system. We have already seen in the previous section that the work done on the system can be calculated as the mismatch between the internal energies of the system and the bath. However, we observe here that by using the canonical Hamiltonian, we can get non-zero work solely due to the time dependence of the system’s Hamiltonian, Eq. (V). This analysis was introduced in [68], where the time dependence of $H^{\rm can}_{S}(t)$ was held responsible for the exchange of energy between the open system and the bath. As a result of this definition of heat exchange, entropy production can be defined by
@@
But to calculate the time-dependent Hamiltonian $\widetilde{H}_{B}(t)$ of the bath, we need to know the map for the evolution of the bath, which is a highly non-trivial task. Nevertheless, we can adopt a different way of calculating the change in the energy of the bath $\Delta\widetilde{U}_{B}(t)$. Analogous to Eq. (IV.1), we can compute the bath heat current in the present scenario. Previously, we conditioned the bath Hamiltonian to be time-independent and derived the expression for the change in the internal energy of the bath. In the present scenario, the rate of change in the internal energy of the bath is given by $\Delta\dot{\widetilde{U}}_{B}=\frac{d}{dt}\left\langle\widetilde{H}_{B}(t)\right\rangle$. Again, using the relation between the expectation of any quantum mechanical operator and the expectation value of the commutator between the operator and the total Hamiltonian of the system, we can rewrite $\Delta\dot{\widetilde{U}}_{B}$ as
@@
The thermodynamic quantities computed in this section using the canonical Hamiltonian are plotted in Fig. 4, along with the corresponding quantities from the previous section to benchmark their variations. From an immediate inspection, we see that the quantities calculated using the canonical Hamiltonian are ambiguous. The entropy production, Fig. 4(c), can be observed to be negative when computed using Eq. (46), violating the second law of thermodynamics. This equation of entropy production has two problems. First, it is calculated in the system’s subspace, albeit considering the canonical Hamiltonian carries the effect of the bath even after it is traced out. A similar observation of this expression of entropy production being negative was also observed in [86]. The other issue is that even if we replace the canonical Hamiltonian with the system’s Hamiltonian and further take it to be time-independent, the expression for the entropy production $\Sigma$ from Eq. (28) doesn’t match Eq. (46) due to the presence of work done by the bath $W$. Furthermore, it is observed from Fig. 4(c) that the entropy production $\Sigma$ in the form of Eq. (28) preserves the second law of thermodynamics ($\Sigma\geq 0$) for the same set of parameters. We also plot the quantity $\Sigma^{\prime}$, Eq. (53), where we replace the internal energy change of the bath $Q_{B}$ in Eq. (28) by the internal energy change of the bath due to time-dependent bath Hamiltonian $\Delta\widetilde{U}_{B}(t)$. However, this quantity also goes negative at various points in time. We refrain from calling it an expression of entropy production because even if we consider a time-dependent bath Hamiltonian, the entropy production [Eq. (27)], for the initial thermal state of the bath, results in Eq. (28) and not in Eq. (53).
@@
The involvement of the canonical Hamiltonian in calculating the heat exchange and work done further poses an issue with the first law of thermodynamics. In this section, we have seen that there are two possible definitions of the first law of thermodynamics. In the first, $\delta\widetilde{W}_{S}(t)$ and $\delta\widetilde{Q}_{S}(t)$ from Eq. (V) can be clubbed together to represent the first law of thermodynamics. Here, non-zero work is obtained because of the canonical Hamiltonian $H^{\rm can}_{S}(t)$. However, from Figs. 4(a) and (b), it can be observed that at the start of the dynamics, the change in the internal energy of the system $\Delta\widetilde{U}_{S}(t)$ is positive, and so is the change in the internal energy of the bath $\Delta\widetilde{U}_{B}(t)$. Further, in the same temporal regime, the work done $\delta\widetilde{W}_{S}$ is negative, Fig. 4(d), depicting that the system works on the bath, and the heat is transferred to the system as $\delta\widetilde{Q}_{S}(t)$ is positive. This observation reveals that on using the canonical Hamiltonian, we get a scenario where the system and the bath both gain energy, and the system works on the bath while heat is transferred to the system. This scenario is ambiguous as it requires some external source to supply energy, which is absent here. The other definition of the first law of thermodynamics comes from Eq. (52). In this scenario, again, we have the same problem where the system and the bath both gain energy at the same time. However, the work done by the bath, which is the mismatch between the system and the bath energies $W=\Delta\widetilde{U}_{S}(t)+\Delta\widetilde{U}_{B}(t)$, is initially positive when $\Delta\widetilde{U}_{S}(t)$ and $\Delta\widetilde{U}_{B}(t)$ are positive. In this case, the work is done by the bath on the system, in contradiction to the work done $\delta\widetilde{W}_{S}$ being negative, as seen above. Therefore, there is an ambiguity in the definition of the work, too. These issues make the formulation of the first law using canonical Hamiltonian ambiguous.
@@ VI Conclusion
This work is devoted to understanding the quantum thermodynamics of strongly coupled non-Markovian quantum systems. In this context, correct definitions of the fundamental quantum thermodynamic quantities, such as heat, work, entropy production, and ergotropy, as well as the laws of quantum thermodynamics, need to be carefully examined. We addressed this issue by taking up a non-trivial model of a central spin immersed in a spin bath and derived its exact reduced dynamics. This was then utilized to calculate the above thermodynamic quantities using the original system and bath Hamiltonian, where the work done by the bath is the deficit between the system and bath internal energies. This led to appropriate definitions of various thermodynamic quantities and provided consistency with the laws of thermodynamics. Thus, for example, a positive entropy production, verifying the second law of quantum thermodynamics, was observed. However, the entropy production rate was sometimes negative, which is typical behavior of non-Markovian quantum systems. An interesting result, relevant to the charging-discharging behavior of the central spin, was obtained using ergotropy. In the present case of the spin bath, we arrived at a situation where the bath itself could act as a charger, such that it gained positive ergotropy even if initially the ergotropy was zero. This is in contrast to what was observed earlier for a harmonic oscillator bath acting as a charger and, hence, puts the role of the central spin model as a quantum battery in perspective. Furthermore, the above thermodynamic quantities were calculated using the recently developed canonical Hamiltonian approach, which led to inconsistencies in the definition of the first and second laws of thermodynamics. It is hoped that this work will provide a stepping stone towards understanding the quantum thermodynamics of strongly coupled non-Markovian systems.
@@ Acknowledgements
Appendix A: The spectral decomposition of the total Hamiltonian
@@
|$\ket{\psi_{\pm}(n)}=\frac{1}{\sqrt{1+\chi_{\pm}(n)^{2}}}\left(\chi_{\pm}\ket{0}_{S}\ket{n}_{B}+\ket{1}_{S}\ket{n-1}_{B}\right),$ (56)

### v2

@@ IV.3 The canonical Hamiltonian of the central spin model
$\Im(\cdot)$ denotes the imaginary part of ()($\cdot$)( ⋅ ). Further, $\Omega(0)=\frac{\omega_{0}}{2}$.
@@
In Fig. 5, we plot the factor $\Omega(t)$ with time $t$. As time progresses, the factor $\Omega(t)$ shows an oscillatory behavior with time around the line of $\frac{\omega_{0}}{2}$; however, it is asymmetric. The minimal dissipator can be found using Eq. (II.3).
@@
IV.3.1 Role of the canonical Hamiltonian in central spin’s thermodynamics
In the preceding section, we discussed the quantum thermodynamics of the central spin system using the system’s bare Hamiltonian $H_{S}$ and the reduced state of the system at any time $t$. However, the dynamics of the reduced state of the system is obtained from a quantum master equation having a canonical Hamiltonian $H^{\rm can}_{S}$, as discussed above. Therefore, it becomes natural to discuss the role of the canonical Hamiltonian in calculating the thermodynamical observables [71, 72]. To this end, we calculate the change in energy of the system by replacing the system’s Hamiltonian with the canonical Hamiltonian, that is,
@@ IV.3.1 Role of the canonical Hamiltonian in central spin’s thermodynamics
One may argue that the canonical Hamiltonian contains the effect of the bath even when the bath is partially traced and that it should be used to calculate expectation values of the thermodynamical quantities. By this argument, we observe that the system’s Hamiltonian becomes time-dependent even if we take a time-independent initial Hamiltonian of the system. We have already seen in the previous section that the work done on the system can be calculated as the mismatch between the internal energies of the system and the bath. However, we observe here that by using the canonical Hamiltonian, we can get non-zero work solely due to the time dependence of the system’s Hamiltonian, Eq. (IV.3.1). This analysis was introduced in [71], where the time dependence of $H^{\rm can}_{S}(t)$ was held responsible for the exchange of energy between the open system and the bath. As a result of this definition of heat exchange, entropy production can be defined by
@@
But to calculate the time-dependent Hamiltonian $\widetilde{H}_{B}(t)$ of the bath, we need to know the map for the evolution of the bath, which is a highly non-trivial task. Nevertheless, we can adopt a different way of calculating the change in the energy of the bath $\Delta\widetilde{U}_{B}(t)$. Analogous to Eq. (II.1), we can compute the bath heat current in the present scenario. Previously, we conditioned the bath Hamiltonian to be time-independent and derived the expression for the change in the internal energy of the bath. In the present scenario, the rate of change in the internal energy of the bath is given by $\Delta\dot{\widetilde{U}}_{B}=\frac{d}{dt}\left\langle\widetilde{H}_{B}(t)\right\rangle$. Again, using the relation between the expectation of any quantum mechanical operator and the expectation value of the commutator between the operator and the total Hamiltonian of the system, we can rewrite $\Delta\dot{\widetilde{U}}_{B}$ as
@@
The thermodynamic quantities computed in this section using the canonical Hamiltonian are plotted in Fig. 6, along with the corresponding quantities from the previous section to benchmark their variations. From an immediate inspection, we see that the quantities calculated using the canonical Hamiltonian show enigmatic behavior. The entropy production, Fig. 6(c), can be observed to be negative when computed using Eq. (54), violating the second law of thermodynamics. This equation of entropy production has two problems. First, it is calculated in the system’s subspace, albeit considering the canonical Hamiltonian carries the effect of the bath even after it is traced out. A similar observation of this expression of entropy production being negative was also observed in [91]. The other issue is that even if we replace the canonical Hamiltonian with the system’s Hamiltonian and further take it to be time-independent, the expression for the entropy production $\Sigma$ from Eq. (6) doesn’t match Eq. (54) due to the presence of work done by the bath $W$. Furthermore, it is observed from Fig. 6(c) that the entropy production $\Sigma$ in the form of Eq. (6) preserves the second law of thermodynamics ($\Sigma\geq 0$) for the same set of parameters. We also plot the quantity $\Sigma^{\prime}$, Eq. (61), where we replace the internal energy change of the bath $Q_{B}$ in Eq. (6) by the internal energy change of the bath due to time-dependent bath Hamiltonian $\Delta\widetilde{U}_{B}(t)$. However, this quantity also goes negative at various points in time. We refrain from calling it an expression of entropy production because even if we consider a time-dependent bath Hamiltonian, the entropy production [Eq. (5)], for the initial thermal state of the bath, results in Eq. (6) and not in Eq. (61).
@@
The involvement of the canonical Hamiltonian in calculating the heat exchange and work done further poses an issue with the first law of thermodynamics. In this section, we have seen that there are two possible definitions of the first law of thermodynamics. In the first, $\delta\widetilde{W}_{S}(t)$ and $\delta\widetilde{Q}_{S}(t)$ from Eq. (IV.3.1) can be clubbed together to represent the first law of thermodynamics. Here, non-zero work is obtained because of the canonical Hamiltonian $H^{\rm can}_{S}(t)$. However, from Figs. 6(a) and (b), it can be observed that at the start of the dynamics, the change in the internal energy of the system $\Delta\widetilde{U}_{S}(t)$ is positive, and so is the change in the internal energy of the bath $\Delta\widetilde{U}_{B}(t)$. Further, in the same temporal regime, the work done $\delta\widetilde{W}_{S}$ is negative, Fig. 6(d), depicting that the system works on the bath, and the heat is transferred to the system as $\delta\widetilde{Q}_{S}(t)$ is positive. This observation reveals that on using the canonical Hamiltonian, we get a scenario where the system and the bath both gain energy, and the system works on the bath while heat is transferred to the system. This scenario is ambiguous as it requires some external source to supply energy, which is absent here. The other definition of the first law of thermodynamics comes from Eq. (60). In this scenario, again, we have the same problem where the system and the bath both gain energy at the same time. However, the work done by the bath, which is the mismatch between the system and the bath energies $W=\Delta\widetilde{U}_{S}(t)+\Delta\widetilde{U}_{B}(t)$, is initially positive when $\Delta\widetilde{U}_{S}(t)$ and $\Delta\widetilde{U}_{B}(t)$ are positive. In this case, the work is done by the bath on the system, in contradiction to the work done $\delta\widetilde{W}_{S}$ being negative, as seen above. Therefore, there is an ambiguity in the definition of the work, too. These issues make the formulation of the first law using the canonical Hamiltonian ill-defined.
@@
V Ergotropy: application of the central spin model as a quantum battery
Here, $H_{S}$ is the system Hamiltonian, and $\rho_{S}^{p}$ is a final state that evolved from a state potent of work production $\rho_{S}(t)$, that is, the passive state. After an external time-dependent cyclic potential has done the maximum work, the state has turned into a passive state, which would be, in general, different from the Gibbs state attained for macroscopic systems. Once the passive state is reached, no more work can be extracted from it through any cyclic process. The spectral decompositions of the passive state and the Hamiltonian of the system are given by
$\rho_{S}^{p}=\sum_{j}r_{j}\ket{\epsilon_{j}}\bra{\epsilon_{j}},$
$H_{S}=\sum_{k}\epsilon_{k}\ket{\epsilon_{k}}\bra{\epsilon_{k}},$ (64)
@@ VI Conclusion
This work is devoted to understanding the quantum thermodynamics of strongly coupled non-Markovian quantum systems. In this context, correct definitions of the fundamental quantum thermodynamic quantities, such as heat, work, entropy production, and ergotropy, as well as the laws of quantum thermodynamics, need to be carefully examined. We addressed this issue by taking up a non-trivial model of a central spin immersed in a spin bath and derived its exact reduced dynamics. This was then utilized to calculate the above thermodynamic quantities using the original system and bath Hamiltonian, where the work done by the bath is the deficit between the system and bath internal energies. This led to appropriate definitions of various thermodynamic quantities and provided consistency with the laws of thermodynamics. Thus, for example, a positive entropy production, verifying the second law of quantum thermodynamics, was observed. However, the entropy production rate was sometimes negative, which is typical behavior of non-Markovian quantum systems. We determined the thermodynamic equilibrium state of the system under strong system-bath coupling using the Hamiltonian of Mean Force (HMF) framework. In the weak coupling and high-temperature regime, the HMF of the system was seen to converge to the bare system Hamiltonian. Additionally, we evaluated the thermodynamic entropy of the system based on the HMF; it was found to be consistent with the third law of thermodynamics. Furthermore, we analyzed the influence of both system-bath coupling strength and temperature on the thermodynamic entropy. Furthermore, the above thermodynamic quantities were also calculated using the recently developed canonical Hamiltonian approach, which led to inconsistencies in the definition of the first and second laws of thermodynamics. An interesting result, relevant to the charging-discharging behavior of the central spin, was obtained using ergotropy. In the present case of the spin bath, we arrived at a situation where the bath itself could act as a charger, such that the system (quantum battery) gained positive ergotropy even if initially the ergotropy was zero. This is in contrast to what was observed earlier for a harmonic oscillator bath acting as a charger and, hence, puts the role of the central spin model as a quantum battery in perspective. It is hoped that this work will provide a stepping stone towards understanding the quantum thermodynamics of strongly coupled non-Markovian systems.
@@ Acknowledgements
Appendix A The spectral decomposition of the total Hamiltonian
@@
|$\ket{\psi_{\pm}(n)}=\frac{\chi_{\pm}(n)\ket{0}_{S}\ket{n}_{B}+\ket{1}_{S}\ket{n-1}_{B}}{\sqrt{1+\chi_{\pm}(n)^{2}}},$ (68)